# Multilingual Mechanism Analysis and Usage with MoE: Scaling Up to 50 Languages in medicine

### v1

@@ 1 Introduction
However, despite the potential of multilingual LLMs in healthcare, significant challenges persist. A major obstacle is the scarcity of medical data in many languages, limiting model development for underrepresented populations. While some research has made strides in addressing this issue, these efforts often focus on only a few dominant languages, typically fewer than six (Wang et al., 2024; Qiu et al., 2024). To address this gap, we have expanded the dataset to include 12 major languages, thereby improving population representation and rigorously evaluating the dataset’s quality and scalability.
Expanding the coverage from 12 major languages to include low-resource languages presents greater challenges due to data scarcity. Addressing this issue requires analyzing the internal information flow of large language models from a multilingual perspective to develop a generalizable approach. While some studies employ neuron analysis to investigate this (Tang et al., 2024), they typically focus on fewer than seven languages, and the complexity of neuron decomposition limits the exploration of relationships between languages and the scalability of applications. To enhance the understanding of these mechanisms, we leverage the modularity of Mixture of Experts models and introduce Hybrid-$k$ routing. This method not only ensures the activation of language-specific experts but also achieves performance on par with the vanilla top-$k$ approach, balancing interpretability and performance.
Inspired by the theory of Circuits (Olah et al., 2020a), we regard experts as nodes, and consider the directed acyclic graph formed by the information transmission of each token from shallow to deep layers as ”Circuits”. Through observation and analysis, we identify the “Spread Out in the End” mechanism in the information flow circuits, where cross-lingual integration occurs in the early layers, while language-specific differentiation happens in the later layers. This insight suggests that employing the Post-MoE architecture, where the MoE structure is applied only in the later layers. Experiments demonstrates that the performance of 12 major languages remained stable, while low-resource languages improved even without additional training.
@@
The main contributions are as follows: 1) We construct a high-quality medical dataset encompassing 12 major languages, quality of which is validated by experiments. 2) We propose a new circuits-based paradigm for interpreting routing in a multilingual context. Through circuit analysis, we identify the “Spread Out in the End” mechanism. 3) By introducing language family experts, we efficiently extend medical LLMs to 50 languages, demonstrating its potential for scaling to more languages.
@@ 2.1 Data Collection and Procession
Evaluation Setup  To ensure the validity of the benchmark, we utilize the multilingual medical benchmark that is publicly available and peer-reviewed. For evaluating languages with limited resources, we follow the multilingual evaluation methodology of Llama3 (Dubey et al., 2024), employing Google Translate to translate the questions and answers of MMLU (Hendrycks et al., 2020). We employed a random selection of 3-shot queries to pose questions to the model, followed by answer extraction and evaluation of the model’s responses. For detailed information regarding the composition of the evaluation set and the evaluation sample used, please refer to the App. A.3.
@@ 2.2 Experimental Results
While we ensure data quality during source evaluation, we also perform a statistical analysis of the dataset’s quality and composition. First, we conduct monolingual training, which involves training the model solely on data from a specific language to assess dataset’s quality in that language. Next, we perform an ablation study focused on the function of code and math to determine the necessity of including this data. Multilingual training, which incorporates all available data using a random sampler, serves as our default reference. We select Gemma-2b (Mesnard et al., 2024) as our base model due to its moderate number of parameters. The evaluation setting and training setting used in this section are consistent with other experiments, see App. B.
@@ 3.1 A Language-specific Hybrid Routing
While Sec. 2 utilizes a dense model to integrate multilingual data and extend the medical model to 12 languages, this approach presents efficiency limitations. To address these challenges, we adopt a sparse Mixture of Experts (MoE) model for a better balance between effectiveness and efficiency. The modular and functional nature of the Experts in the MoE framework is particularly advantageous in further scalability, which benefit enabling the effective expansion to more languages (especially for minor languages in Sec. 4.
@@ 3.1.2 Experimental Comparisons between Routing Strategies
Experiment Settings Considering its broad parameter base and impressive multilingual capabilities, we selected the Qwen2 series for our experiments. Specifically, we conducted experiments using the Qwen2-0.5B model, based on the training and evaluation datasets described in Sec. 2. For token language classification, tokens are uniformly classified based on their source per document. To construct the evaluation set for minor languages, we evenly select 38 languages based on their geographical distribution. Similar to Sec. 2 , we use Google Translate to translate the questions and answers from the medical-clinical section of the MMLU dataset, types of minor languages are detailed in the App. A.5. For Lang-Spec. routing, we use 12 experts plus one shared expert. For Hybrid-$k$ and Top-$k$ routing, 12 experts are used per layer. The activation count is fixed at two experts across all configurations ($k=2$) if not specified. Additional training settings are provided in App. B.
@@
Method Avg. Accuracy Ar De En Es Fr Hi It Ja Ko Pt Ru Zh
Major Minor
Dense Models before and after Training
Qwen-0.5B 29.7 31.5 27.3 27.4 39.3 32.9 21.3 25.7 20.7 24.0 19.2 26.3 46.9 45.4
after training 37.8 24.6 34.8 33.7 46.7 40.5 33.0 31.0 31.5 31.5 35.5 30.8 53.8 51.0
@@
Lang-Spec. 30.9 26.1 28.8 28.6 39.1 33.3 16.5 24.5 21.4 27.8 31.0 52.3 42.8 53.0
Top-$k$ 39.7 29.9 34.5 36.9 43.7 38.9 39.9 32.2 37.7 30.5 35.9 34.8 58.2 53.2
Hybrid-$k$ 40.0 32.0 35.1 37.2 44.1 40.8 40.7 28.7 38.9 32.4 36.2 34.3 58.8 53.6
Results  Tab. 2 shows that the base model in major languages is improved significantly after fine-tuning, but its performance in minor languages declines substantially. This indicates that partial language fine-tuning of dense models notably impacts their generalization capabilities to other languages. The Language-Specific routing of MoE provides interpretability but results in poor multilingual performance. In contrast, the Top-$k$ MoE model demonstrates superior multilingual capability and generalization after fine-tuning compared to the trained base model, highlighting the effectiveness of MoE models over dense models. Moreover, Hybrid-$k$ routing exhibits a clear advantages than Top-$k$ in nearly all languages, especially in minor languages; this evidences its superior generalization.
@@ 3.2.1 Formulation of Information Flow in Routing
Definition 1. Information Flow Circuit The sequence of Experts that each token passes from shallow to deeper abstractions forms a directed acyclic graph (DAG), which we refer to as a ‘circuit’.
@@ 3.2.2 Results
The visualization of the routing pattern is shown in Fig. 3. From the perspective of single-language routing, the routing results for Japanese and Spanish in the figure (routing distributions for other languages are detailed in the App.E) indicate that Chinese has a significant influence on Japanese, while Portuguese and English are also integrated into the information flow of Spanish. This phenomenon aligns well with linguistic priors: the linguistic development of Japanese has been influenced by Chinese, while Spanish, Portuguese, and English, all belonging to the Romance language family, have historically exerted mutual influence. From the perspective of cross-linguistic routing for all 12 languages, through the utilization of routing distribution analysis and Information Flow Circuits visualization, we observe that the information flow circuits exhibit cross-linguistic concentration in the early layers, while differentiation based on language occurs in the later layers. We refer to this phenomenon as “Spread Out in the End”:
Phenomenon 1. In the early layers, the model exhibits shared routing patterns across multiple languages. However, in the later layers, the model diverges, with tokens being routed to language-specific experts, allowing late routing to specialize in its respective language.
@@ 3.3 A MoE variant inspired by the ‘Spread Out in the End’ phenomenon: Post-MoE
Figure 4: Analysis of Upcycling Layer Depths for the PostMoE Architecture. The X-axis represents the number of Upcycling layers applied in the final N layers, while the Y-axis indicates the model performance on both Major and Minor languages. N=0 signifies direct fine-tuning of the model. Qwen2-0.5B-MoE and Qwen2-1.5B-MoE refer to standard MoE architectures trained with Hybrid routing.
@@
Experiment Settings To validate the effectiveness of the multilingual mechanism and the Post-MoE architecture, we use Qwen2-0.5B-MoE and Qwen2-1.5B-MoE, along with the original base model as a baseline. Qwen2-0.5B-MoE and Qwen2-1.5B-MoE are standard MoE architectures trained with Hybrid routing. To investigate the impact of the number of MoE layers in the final layer of the model on performance, we extend the MoE architecture in the last 1, 2, 3, and 4 layers using the hybrid routing method. We also evaluate the multilingual generalization capability of this architecture using the assessment set of 38 minor languages mentioned in Sec. 4.
@@ 4 Further Scaling to 50 Languages
To further demonstrate the multilingual capabilities of the Post-MoE architecture, we selected 38 minor languages, expanding the language variety to 50 (shown in Tab. 3).
@@ 4.1 Mixture of Language Family Experts
Following the Hybrid-$k$ which adopts language-specifc experts, training LLMs with $n$ languages would require $n$ language-specific experts. This expansion strategy would lead to an substantial growth in model parameters, making the application impractical. To address this challenge, we propose adapting Post-MoE by introducing the Mixture of Language Family Experts, which groups the 50 languages into 7 established linguistic families, as detailed in Tab. 3. This approach named ‘Apollo-MoE’ ensures that scaling to additional languages does not necessitate a corresponding increase in parameters. Each language family employs hybrid routing tailored to its linguistic characteristics, enabling more efficient training. This method might facilitate scalable and robust multilingual performance across a wide range of languages.
@@ 4.2 Experiments
Model #params Average Acc. Ar De En Es Fr Hi It Ja Ko Pt Ru Zh
Major Minor
@@
JSL-MedPhi2 2.7B 29.0 30.7 25.0 30.8 42.1 34.6 30.6 13.2 26.6 17.9 21.3 29.7 48.8 26.8
MMed-Llama-3 8B 40.2 36.5 29.5 40.3 54.7 63.6 62.0 38.8 20.7 47.3 20.3 25.9 62.5 16.7
OpenBioLLM 8B 46.1 34.7 27.5 58.1 49.6 59.2 53.3 44.4 41.5 39.1 46.7 27.0 64.8 41.8
Llama3 8B 49.3 33.3 33.5 55.9 48.4 60.5 58.4 48.0 39.9 38.9 49.5 51.5 63.3 43.6
@@
Apollo-MoE 0.7B 40.5 34.6 36.3 38.2 45.4 39.8 38.4 33.1 39.9 26.9 35.2 37.3 64.1 51.3
Apollo-MoE 2.0B 54.8 44.9 47.2 53.8 56.5 52.5 53.3 39.5 54.4 45.7 49.5 57.3 69.1 66.8
Apollo-MoE 10.1B 69.9 58.3 58.3 73.5 73.1 69.4 72.4 56.9 71.9 62.4 68.4 73.8 74.2 84.1
@@
Experiment Settings  Given the extremely limited data for these rare languages, we extracted 2,000 samples from English data and used Google Translate to create training corpora for each language. The clinical-knowledge section of MMLU was translated to serve as the corresponding evaluation set. We trained the Post-MoE model, named Apollo-MoE, using Qwen2-0.5B, 1.5B, and 7B as base models with a Mixture of Language Family Experts. We evaluated Apollo-MoE on benchmarks for major and minor languages and compared it with open-source and high-performing open-source medical models.
@@
Results  Tab. 4 demonstrate that our Apollo-MoE models outperform other models of similar sizes in both major and minor languages. The 2B model achieves 54.8 in major and 44.9 in minor languages, surpassing open-source models with 8B parameters. The 10B model leads all open-source models, achieving 69.9 in major and 58.3 in minor languages. Notably, Apollo-MoE excels in major languages like English, French, and Spanish, with particularly strong performance in French, and outperforms other models significantly in low-resource languages like Arabic and Hindi.
@@
Ablation Study  Tab. 5 provides a detailed comparison of routing performance across 50 languages, with two experts routed among 7 linguistic family experts in the Post-MoE model. It shows that routing strategies consistently outperform Dense models, with Hybrid-$k$ routing achieving an average major language accuracy of 54.8 for Qwen2-1.5B, compared to 52.2 for Dense. Additionally, Hybrid-$k$ routing shows a clear advantage over Top-$k$ routing, particularly in minor languages, where the Qwen2-7B model achieves an average accuracy of 58.3 with Hybrid-$k$, compared to 56.7 with Top-$k$.
Table 5: Ablation Study on 50 Languages.
Languages Average acc. Ha Sr La Gn Doi Da Ln Ceb Mai Mg Ilo
Major Minor
Qwen2-0.5B 29.7 31.5 31.1 30.7 36.4 28.4 27.7 33.3 29.9 34.5 27.3 34.8 30.7
Dense 39.2 33.2 33.7 25.0 30.7 33.3 32.6 36.0 34.5 36.7 31.8 34.1 35.2
Top-$k$ routing 39.0 33.4 26.9 17.8 37.5 34.1 29.9 36.7 33.7 32.2 28.8 32.2 32.6
Hybrid-$k$ routing 40.5 34.6 31.1 25.4 38.3 34.8 36.0 39.8 41.3 40.9 33.7 37.9 39.8
Qwen2-1.5B 42.9 38.4 36.0 44.7 39.0 37.5 31.8 45.1 34.8 42.4 34.1 39.0 35.6
Dense 52.2 43.7 36.7 30.3 51.1 47.0 37.1 51.1 44.3 48.9 39.8 40.5 45.8
Top-$k$ routing 53.6 42.6 39.8 24.6 47.7 44.7 36.0 50.4 34.5 48.5 35.6 41.7 43.9
Hybrid-$k$ routing 54.8 44.9 39.0 33.7 48.5 45.8 41.3 55.3 43.9 50.8 42.8 41.7 46.2
Qwen2-7B 55.2 49.2 34.1 57.6 52.3 43.2 40.9 63.3 38.3 58.3 48.5 37.1 45.5
Dense 69.0 55.7 37.9 31.1 60.6 54.9 55.7 68.6 47.3 64.4 55.7 44.7 64.4
Top-$k$ Routing 68.6 56.7 37.1 39.4 60.6 56.1 50.8 67.8 53.4 67.0 59.8 46.2 61.0
Hybrid-$k$ Routing 69.9 58.3 48.5 42.0 62.9 61.0 54.5 71.2 53.0 68.6 58.3 50.8 64.4
Figure 5: Data Scale ACC.
@@
On the Data Scale  To investigate the sensitivity to data scale, we trained Post-MoE model with Qwen2-0.5B base using a combination of major languages along with 10%, 30%, 50%, and 100% of the data from minor languages, corresponding to 200, 600, 1,000, and 2,000 data examples per language, respectively. As shown in Fig. 5, the model’s performance across various languages improves with increasing data but eventually plateaus. This indicates that the proposed method is not heavily data-dependent, reaching saturation with as few as 2,000 data examples; this is especially beneficial for low-resources languages.

### v2

# Efficiently Democratizing Medical LLMs for 50 Languages via a Mixture of Language Family Experts

@@ 1 Introduction
However, despite the potential of multilingual LLMs in healthcare, significant challenges persist. A major obstacle is the scarcity of medical data in many languages, limiting model development for underrepresented populations. While some research has made strides in addressing this issue, these efforts often focus on only a few dominant languages, typically fewer than six (Wang et al., 2024; Qiu et al., 2024). To address this gap, we have expanded the dataset to include 12 high-resource languages, thereby improving population representation and rigorously evaluating the dataset’s quality and scalability.
Expanding the coverage from 12 high-resource languages to include low-resource languages presents greater challenges due to data scarcity. Addressing this issue requires analyzing the internal information flow of large language models from a multilingual perspective to develop a generalizable approach. While some studies employ neuron analysis to investigate this (Tang et al., 2024), they typically focus on fewer than seven languages, and the complexity of neuron decomposition limits the exploration of relationships between languages and the scalability of applications. To enhance the understanding of these mechanisms, we leverage the modularity of Mixture of Experts models and introduce Hybrid-$k$ routing. This method not only ensures the activation of language-specific experts but also achieves performance on par with the vanilla top-$k$ approach, balancing interpretability and performance.
Inspired by the theory of Circuits (Olah et al., 2020a), we regard experts as nodes, and consider the directed acyclic graph formed by the information transmission of each token from shallow to deep layers as ”Circuits”. Through observation and analysis, we identify the “Spread Out in the End” mechanism in the information flow circuits, where cross-lingual integration occurs in the early layers, while language-specific differentiation happens in the later layers. This insight suggests that employing the Post-MoE architecture, where the MoE structure is applied only in the later layers. Experiments demonstrates that the performance of 12 high-resource languages remained stable, while low-resource languages improved even without additional training.
@@
The main contributions are as follows: 1) We construct a high-quality medical dataset encompassing 12 high-resource languages, quality of which is validated by experiments. 2) We propose a new circuits-based paradigm for interpreting routing in a multilingual context. Through circuit analysis, we identify the “Spread Out in the End” mechanism. 3) By introducing language family experts, we efficiently extend medical LLMs to 50 languages, demonstrating its potential for scaling to more languages.
@@ 2.1 Data Collection and Procession
Evaluation Setup  To ensure the validity of the benchmark, we utilize the multilingual medical benchmark that is publicly available and peer-reviewed. For evaluating languages with limited resources, we follow the multilingual evaluation methodology of Llama3 (Dubey et al., 2024), employing Google Translate to translate the questions and answers of MMLU (Hendrycks et al., 2020). We also further verified the effectiveness of Google Translate in related aspects through experiments detailed in App.A.6. We employed a random selection of 3-shot queries to pose questions to the model, followed by answer extraction and evaluation of the model’s responses. For detailed information regarding the composition of the evaluation set and the evaluation sample used, please refer to the App. A.3.
@@ 2.2 Experimental Results
While we ensure data quality during source evaluation, we also perform a statistical analysis of the dataset’s quality and composition. First, we conduct monolingual training, which involves training the model solely on data from a specific language to assess dataset’s quality in that language. Next, we perform an ablation study focused on the function of code and math to determine the necessity of including this data. Multilingual training, which incorporates all available data using a random sampler, serves as our default reference. We select Gemma-2b (Team et al., 2024a) as our base model due to its moderate number of parameters. The evaluation setting and training setting used in this section are consistent with other experiments, see App. B.
@@ 3.1 A Language-specific Hybrid Routing
While Sec. 2 utilizes a dense model to integrate multilingual data and extend the medical model to 12 languages, this approach presents efficiency limitations. To address these challenges, we adopt a sparse Mixture of Experts (MoE) model for a better balance between effectiveness and efficiency. The modular and functional nature of the Experts in the MoE framework is particularly advantageous in further scalability, which benefit enabling the effective expansion to more languages (especially for low-resource languages in Sec. 4.
@@ 3.1.2 Experimental Comparisons between Routing Strategies
Experiment Settings Considering its broad parameter base and impressive multilingual capabilities, we selected the Qwen2 series for our experiments. Specifically, we conducted experiments using the Qwen2-0.5B model, based on the training and evaluation datasets described in Sec. 2. To accurately construct a dense model of equivalent size with corresponding MoE model as baseline, we replicated the MLP following the approach used in MoE Upcycling and initialized the routing with an average distribution. Unlike MoE, the initialized dense model employs full activation instead of sparse activation. For token language classification, tokens are uniformly classified based on their source per document. To construct the evaluation set for low-resource languages, we evenly select 38 languages based on their geographical distribution. Similar to Sec. 2, we use Google Translate to translate the questions and answers from the medical-clinical section of the MMLU dataset, types of low-resource languages are detailed in the App. A.5. For Lang-Spec. routing, we use 12 experts plus one shared expert. For Hybrid-$k$ and Top-$k$ routing, 12 experts are used per layer. The activation count is fixed at two experts across all configurations ($k=2$) if not specified. Additional training settings are provided in App. B.
@@
Method Param. (B) Avg. Accuracy Ar De En Es Fr Hi It Ja Ko Pt Ru Zh
Active Total High Low
Dense Models before and after Training with Various Param.
Qwen-0.5B 0.49 0.49 29.7 31.5 27.3 27.4 39.3 32.9 21.3 25.7 20.7 24.0 19.2 26.3 46.9 45.4
after training 0.49 0.49 37.8 24.6 34.8 33.7 46.7 40.5 33.0 31.0 31.5 31.5 35.5 30.8 53.8 51.0
Same Active 0.81 0.81 38.4 26.2 34.4 35.8 46.1 40.1 34.0 32.2 32.2 32.0 35.1 31.0 54.5 51.8
Same Total 3.95 3.95 42.0 30.9 36.4 40.8 47.8 42.1 38.0 34.6 43.2 32.4 38.7 31.9 62.9 54.8
@@
Lang-Spec. 0.81 3.95 30.9 26.1 28.8 28.6 39.1 33.3 16.5 24.5 21.4 27.8 31.0 52.3 42.8 53.0
Top-$k$ 0.81 3.95 39.7 29.9 34.5 36.9 43.7 38.9 39.9 32.2 37.7 30.5 35.9 34.8 58.2 53.2
Hybrid-$k$ 0.81 3.95 40.0 32.0 35.1 37.2 44.1 40.8 40.7 28.7 38.9 32.4 36.2 34.3 58.8 53.6
Results  Tab. 2 shows that the base model in high-resource languages is improved significantly after fine-tuning, but its performance in low-resource languages declines substantially. This indicates that partial language fine-tuning of dense models notably impacts their generalization capabilities to other languages. The Language-Specific routing of MoE provides interpretability but results in poor multilingual performance. In contrast, the Top-$k$ MoE model demonstrates superior multilingual capability and generalization after fine-tuning compared to the trained base model, highlighting the effectiveness of MoE models over dense models. Moreover, Hybrid-$k$ routing exhibits a clear advantages than Top-$k$ in nearly all languages, especially in low-resource languages; this evidences its superior generalization.
@@ 3.2.1 Formulation of Information Flow in Routing
Definition 1.
Information Flow Circuit The sequence of Experts that each token passes from shallow to deeper abstractions forms a directed acyclic graph (DAG), which we refer to as a ‘circuit’.
@@ 3.2.2 Results
The visualization of the routing pattern is shown in Fig. 3. From the perspective of single-language routing, the routing results for Japanese and Spanish in the figure (routing distributions for other languages are detailed in the App.G) indicate that Chinese has a significant influence on Japanese, while Portuguese and English are also integrated into the information flow of Spanish. This phenomenon aligns well with linguistic priors: the linguistic development of Japanese has been influenced by Chinese, while Spanish, Portuguese, and English, all belonging to the Romance language family, have historically exerted mutual influence. From the perspective of cross-linguistic routing for all 12 languages, through the utilization of routing distribution analysis and Information Flow Circuits visualization, we observe that the information flow circuits exhibit cross-linguistic concentration in the early layers, while differentiation based on language occurs in the later layers. We refer to this phenomenon as “Spread Out in the End”:
Phenomenon 1.
In the early layers, the model exhibits shared routing patterns across multiple languages. However, in the later layers, the model diverges, with tokens being routed to language-specific experts, allowing late routing to specialize in its respective language.
@@ 3.3 A MoE variant inspired by the ‘Spread Out in the End’ phenomenon: Post-MoE
Figure 4: Analysis of Upcycling Layer Depths for the PostMoE Architecture. The X-axis represents the number of Upcycling layers applied in the final N layers, while the Y-axis indicates the model performance on both high- and low-resource languages. N=0 signifies direct fine-tuning of the model. Qwen2-0.5B-MoE and Qwen2-1.5B-MoE refer to standard MoE architectures trained with Hybrid routing.
@@
Experiment Settings To validate the effectiveness of the multilingual mechanism and the Post-MoE architecture, we use Qwen2-0.5B-MoE and Qwen2-1.5B-MoE, along with the original base model as a baseline. Qwen2-0.5B-MoE and Qwen2-1.5B-MoE are standard MoE architectures trained with Hybrid routing. To investigate the impact of the number of MoE layers in the final layer of the model on performance, we extend the MoE architecture in the last 1, 2, 3, and 4 layers using the hybrid routing method. We also evaluate the multilingual generalization capability of this architecture using the assessment set of 38 low-resource languages mentioned in Sec. 4.
@@ 4 Further Scaling to 50 Languages
To further demonstrate the multilingual capabilities of the Post-MoE architecture, we selected 38 low-resource languages, expanding the language variety to 50 (shown in Tab. 3).
@@ 4.1 Mixture of Language Family Experts
Following the Hybrid-$k$ which adopts language-specifc experts, training LLMs with $n$ languages would require $n$ language-specific experts. This expansion strategy would lead to an substantial growth in model parameters, making the application impractical. To address this challenge, we propose adapting Post-MoE by introducing the Mixture of Language Family Experts, which groups the 50 languages into 7 established linguistic families, as detailed in Tab. 3. This approach named ‘Apollo-MoE’ ensures that scaling to additional languages does not necessitate a corresponding increase in parameters. Each language family employs hybrid routing tailored to its linguistic characteristics, enabling more efficient training. This method might facilitate scalable and robust multilingual performance across a wide range of languages. After training 50 languages on Dense models and ablation study of various routing strategies in PostMoE, we further verify the effective multilingual generalization of Apollo-MoE in App.E.
@@ 4.2 Experiments
Model Active Average Acc. Ar De En Es Fr Hi It Ja Ko Pt Ru Zh
Param. High Low
@@
JSL-MedPhi2 2.78 B 29.0 30.7 25.0 30.8 42.1 34.6 30.6 13.2 26.6 17.9 21.3 29.7 48.8 26.8
MMed-Llama-3 8.03 B 40.2 36.5 29.5 40.3 54.7 63.6 62.0 38.8 20.7 47.3 20.3 25.9 62.5 16.7
OpenBioLLM 8.03 B 46.1 34.7 27.5 58.1 49.6 59.2 53.3 44.4 41.5 39.1 46.7 27.0 64.8 41.8
Llama3 8.03 B 49.3 33.3 33.5 55.9 48.4 60.5 58.4 48.0 39.9 38.9 49.5 51.5 63.3 43.6
@@
Apollo-MoE 0.52 B 40.5 34.6 36.3 38.2 45.4 39.8 38.4 33.1 39.9 26.9 35.2 37.3 64.1 51.3
Apollo-MoE 1.63 B 54.8 44.9 47.2 53.8 56.5 52.5 53.3 39.5 54.4 45.7 49.5 57.3 69.1 66.8
Apollo-MoE 8.02 B 69.9 58.3 58.3 73.5 73.1 69.4 72.4 56.9 71.9 62.4 68.4 73.8 74.2 84.1
@@
Experiment Settings  Given the extremely limited data for these rare languages, we extracted 2,000 samples from English data and used Google Translate to create training corpora for each language. The clinical-knowledge section of MMLU was translated to serve as the corresponding evaluation set. We trained the Post-MoE model, named Apollo-MoE, using Qwen2-0.5B, 1.5B, and 7B as base models with a Mixture of Language Family Experts. We evaluated Apollo-MoE on benchmarks for high-resource (see App. 7) and low-resource languages and compared it with close-source and high-performing open-source medical models.
Figure 5: Data Scale Performance.
@@
Results Tab. 4 demonstrate that our Apollo-MoE models outperform other models of similar sizes in both high- and low-resource languages. The 2B model (1.63B active) achieves 54.8 in high-resource and 44.9 in low-resource languages, surpassing open-source models with 8B parameters. The 10B model (8.02B active) leads all open-source models, achieving 69.9 in high-resource and 58.3 in low-resource languages. Notably, Apollo-MoE excels in high-resource languages like English, French, and Spanish, with particularly strong performance in French, and outperforms other models significantly in low-resource languages like Arabic and Hindi.
@@
Ablation Study  Tab. 5 provides a detailed comparison of routing performance across 50 languages, with two experts routed among 7 linguistic family experts in the Post-MoE model (k=2). It shows that routing strategies consistently outperform Dense models, with Hybrid-$k$ routing achieving an average high-resource language accuracy of 54.8 for Qwen2-1.5B, compared to 52.2 for Dense. Additionally, Hybrid-$k$ routing shows a clear advantage over Top-$k$ routing, particularly in low-resource languages, where the Qwen2-7B model achieves an average accuracy of 58.3 with Hybrid-$k$, compared to 56.7 with Top-$k$.
@@
On the Data Scale  To investigate the sensitivity to data scale, we trained Post-MoE model with Qwen2-0.5B base using a combination of high-resource languages along with 10%, 30%, 50%, and 100% of the data from low-resource languages, corresponding to 200, 600, 1,000, and 2,000 data examples per language, respectively. As shown in Fig. 5, the model’s performance across various languages improves with increasing data but eventually plateaus. This indicates that the proposed method is not heavily data-dependent, reaching saturation with as few as 2,000 data examples; this is especially beneficial for low-resources languages.
Table 5: Ablation Study between Dense and PostMoE Models across 50 Languages at Various Scales.
Languages Active Average acc. Ha Sr La Gn Doi Da Ln Ceb Mai Mg Ilo
Param. High Low
Qwen2-0.5B 0.49 B 29.7 31.5 31.1 30.7 36.4 28.4 27.7 33.3 29.9 34.5 27.3 34.8 30.7
Dense 0.49 B 39.2 33.2 33.7 25.0 30.7 33.3 32.6 36.0 34.5 36.7 31.8 34.1 35.2
Dense Same Active 0.52 B 39.4 34.0 34.1 25.5 32.7 34.2 34.2 34.4 38.9 39.2 32.5 32.2 35.9
Top-$k$ routing 0.52 B 39.0 33.4 26.9 17.8 37.5 34.1 29.9 36.7 33.7 32.2 28.8 32.2 32.6
Hybrid-$k$ routing 0.52 B 40.5 34.6 31.1 25.4 38.3 34.8 36.0 39.8 41.3 40.9 33.7 37.9 39.8
Qwen2-1.5B 1.54 B 42.9 38.4 36.0 44.7 39.0 37.5 31.8 45.1 34.8 42.4 34.1 39.0 35.6
Dense 1.54 B 52.2 43.7 36.7 30.3 51.1 47.0 37.1 51.1 44.3 48.9 39.8 40.5 45.8
Dense Same Active 1.63 B 52.8 44.1 38.8 28.3 50.0 49.1 39.9 53.1 44.7 48.1 40.5 40.2 46.5
Top-$k$ routing 1.63 B 53.6 42.6 39.8 24.6 47.7 44.7 36.0 50.4 34.5 48.5 35.6 41.7 43.9
Hybrid-$k$ routing 1.63 B 54.8 44.9 39.0 33.7 48.5 45.8 41.3 55.3 43.9 50.8 42.8 41.7 46.2
Qwen2-7B 7.62 B 55.2 49.2 34.1 57.6 52.3 43.2 40.9 63.3 38.3 58.3 48.5 37.1 45.5
Dense 7.62 B 69.0 55.7 37.9 31.1 60.6 54.9 55.7 68.6 47.3 64.4 55.7 44.7 64.4
Dense Same Active 8.02 B 68.5 56.3 42.2 39.7 59.5 57.9 55.5 69.3 51.1 65.9 57.8 46.8 61.7
Top-$k$ Routing 8.02 B 68.6 56.7 37.1 39.4 60.6 56.1 50.8 67.8 53.4 67.0 59.8 46.2 61.0
Hybrid-$k$ Routing 8.02 B 69.9 58.3 48.5 42.0 62.9 61.0 54.5 71.2 53.0 68.6 58.3 50.8 64.4